Question 1

The most recent financial statements for Heine, Inc., are shown here:

Income Statement Balance Sheet
  Sales $ 23,700   Assets $ 55,200   Debt $ 20,400
  Costs 14,400   Equity 34,800
  Taxable income $ 9,300   Total $ 55,200   Total $ 55,200
  Taxes (40%) 3,720
  
   Net income $ 5,580
  

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $1,800 was paid, and the company wishes to maintain a constant payout ratio. Next year’s sales are projected to be $29,625.

What is the external financing needed? (Do not round intermediate calculations.)

  

  External financing needed $   

Question 2

Al's Sport Store has sales of $2,830, costs of goods sold of $2,120, inventory of $507, and accounts receivable of $419. How many days, on average, does it take the firm to sell its inventory assuming that all sales are on credit?

Multiple Choice

  • 119.3

  • 65.4

  • 87.3

  • 86.1

  • 119.5

Answer ____________
Answer #1

1.Dividend payout ratio=Dividend/Net income

=(1800/5580)=0.322580645

Growth rate in sales=(29625-23700)/23700

=25%

Sales 29625
Costs(14400*1.25) 18000
Taxable income 11625
Tax@40% 4650
Net income $6975
Less:dividend(6975*0.322580645) ($2250)
Addition to retained earnings $4725

Total assets would be=$55200*1.25=$69000

Total equity=$34800+Addition to retained earnings

=$34800+$4725=$39525

Total assets =Total equity+Total liabilities

Hence external financing needed=$69000-39525-$20400

=$9075

b.

Days on average required by firm to sell its inventory=(inventory/Cost of goods sold)*365 days

=(507/2120)*365

=87.3 days(Approx).